OpenStudy (anonymous):

(A.dell)R=A plzzzzz prove it... where A is any costant vector and R=xi+yj+zk

OpenStudy (experimentx):

Did you mean \[ (A * \nabla)r\]

OpenStudy (anonymous):

replace r with R

OpenStudy (anonymous):

(A dot dell)R

OpenStudy (experimentx):

Let \[ A = a_x i + a_y j + a_z k \\ A* \nabla = (a_x i + a_y j + a_z k)* \left ( i\frac \partial {\partial x} + j\frac \partial {\partial y} + k\frac \partial {\partial z}\right ) \\ = a_x\frac \partial {\partial x} + a_y\frac \partial {\partial y} + a_z\frac \partial {\partial z}\] Now we have \[ \left ( a_x\frac \partial {\partial x} + a_y\frac \partial {\partial y} + a_z\frac \partial {\partial z} \right ) ( i x + jy + kz) = i(a_x + 0 + 0) + j(0 + a_y + 0) + k(0 + 0 + a_z) \\ = ia_x + ja_y + ka_z = A\]

OpenStudy (anonymous):

if A is constant delta* A=0

OpenStudy (experimentx):

I guess it was \[ \vec A * \nabla \text{ not } \nabla*\vec A\] sure it would have been true then

OpenStudy (anonymous):

ur right @experimentX

OpenStudy (anonymous):

it's the same thing

OpenStudy (anonymous):

i think A cant be a constant

OpenStudy (experimentx):

Well, i meant the nabla is before or after A matters though i'm not so sure ...

OpenStudy (anonymous):

dot product is commutative

OpenStudy (experimentx):

isn't it supposed to be operator?? i'm not sure though ...

OpenStudy (anonymous):

if oyu have A=(3,3,3),you will get (3,3,3)=(0,0,0)

OpenStudy (experimentx):

lol ... now i'm confused!!

OpenStudy (anonymous):

ax,must depend x,ay y, az z do you agree?

OpenStudy (experimentx):

yes

OpenStudy (experimentx):

but isn't \[ \nabla *(ix+jy+kz) = 3\] while \[ (ix+jy+kz)*\nabla = \text { some operator }\] I think you might encounter them a lot in quantum mechanics